OpenStudy (anonymous):

solve the equation: 2cos^2(x) + 3cos(x) +cos^2(x) + sin^2(x) = 0 where 0 <= x < 2pi

OpenStudy (anonymous):

\[ 2\cos^2(x) + 3\cos(x) +\cos^2(x) + \sin^2(x) =0\\ 2\cos^2(x) + 3\cos(x) +1 =0\\ (\cos(x) +1)( 2\cos(x) +1)=0 \] Can you finish it now?

OpenStudy (anonymous):

\[ \cos(x) =-1\implies x = \pi \text{ Since we need answers between 0 and } 2\pi \]

OpenStudy (anonymous):

\[ \cos(x)=-\frac 1 2\implies x=\frac{2 \pi }3 \text {or } x=\frac{4 \pi }3 \text {Since we need answers between 0 and} 2\pi \]

OpenStudy (anonymous):

whoa never thought to factor XD thank you so much!

OpenStudy (anonymous):

YW